### v2

# Approximation and FPT Algorithms for Finding DM-Irreducible Spanning Subgraphs

Abstract
Finding a minimum-weight strongly connected spanning subgraph of an edge-weighted directed graph is equivalent to the weighted version of the well-known strong connectivity augmentation problem, and it is NP-hard. For this problem, a simple $2$-approximation algorithm was proposed by Frederickson and Jájá (1981); surprisingly, it still achieves the best-known approximation ratio in general. Also, the unweighted problem was shown to be FPT (fixed-parameter tractable) by Bang-Jensen and Yeo (2008), if it is parameterized by the difference from the trivial upper bound of the optimal value. In this paper, we consider a generalized problem related to the Dulmage–Mendelsohn decompositions of bipartite graphs instead of the strong connectivity of directed graphs, and extend these approximation and FPT results to the generalized setting.
Keywords
Connectivity augmentation, Bipartite matching, Matroid intersection, Approximation algorithm, Fixed-parameter tractability.
1 Introduction
A directed graph is called strongly connected if there exists a path between any pair of vertices. The strong connectivity augmentation problem (SCA) is, given a directed graph and the weights of additional edges, to find a minimum-weight subset of additional edges making the graph strongly connected. While SCA is NP-hard in general, it can be solved in linear time in the unweighted case (when any edge can be added and has the same weight) [9].
The following problem, called the strongly connected spanning subgraph problem (SCSS), is polynomial-time equivalent to SCA in general: given a strongly connected directed graph with edge weight, find a minimum-weight spanning subgraph that is strongly connected. Frederickson and Jájá [14] gave a $2$-approximation algorithm for SCSS; this is based on a very simple observation (see Section 3.1) and, surprisingly, it still achieves the best known approximation ratio in general. For the unweighted case (when all the edges have the same weight, which is already NP-hard in contrast to SCA), several algorithms with better approximation ratios [18, 26, 27, 17] have been proposed. Also, Bang-Jensen and Yeo [2] showed that the unweighted case is FPT if it is parameterized by the difference from the trivial upper bound of the optimal value (see Sections 2.2 and 4.1); here, the problem is called FPT (fixed-parameter tractable) if it can be solved in $f(k)\cdot n^{O(1)}$ time, where $k$ and $f$ are the prescribed parameter and computable function, respectively, and $n$ is the input size.
SCSS (or SCA) is generalized to several directions. One natural generalization is the $k$(-edge)-connected spanning subgraph (or $k$(-edge)-connectivity augmentation) problem, where a directed graph is called $k$-connected (or $k$-edge-connected) if there exists $k$ openly disjoint (or edge-disjoint) paths between any pair of vertices. The case when $k=1$ corresponds to SCSS (or SCA), and the general case is also well-studied [10, 12, 19, 24, 1].
Bérczi, Iwata, Kato, and Yamaguchi [3] recently investigated another generalization of SCA, which is related to the Dulmage–Mendelsohn decomposition (the DM-decomposition) [7, 6] of a bipartite graph. The DM-decomposition gives a partition of the vertex set that reflects the structure of all maximum matchings therein, and a bipartite graph is DM-irreducible if its DM-decomposition consists of a single component. This property is closed under adding edges, and we analogously formulate the DM-irreducible spanning subgraph problem (DMISS): given a DM-irreducible bipartite graph with edge weight, find a minimum-weight spanning subgraph that is DM-irreducible. Then, DMISS indeed generalizes SCSS (cf. [3, Section 2.3]), and in terms of DMISS, Bérczi, et al. [3] proposed polynomial-time algorithms for the case when the input bipartite graph is unbalanced with the aid of weighted matroid intersection, and for the case when the input graph is a balanced complete bipartite graph and the weight of each edge is 00 or $1$, where a bipartite graph is balanced if the two color classes have the same number of vertices. In this paper, by extending the approaches to SCSS by Frederickson and Jájá [14] and by Bang-Jensen and Yeo [2], we propose a $2$-approximation algorithm for DMISS in general and an FPT algorithm for the unweighted case of DMISS.
The rest of this paper is organized as follows. In Section 2, we describe necessary definitions, and formally state the problems and our results. In Section 3, we give a $2$-approximation algorithm for DMISS. In Section 4, we prove the unweighted case of DMISS is FPT parameterized by the difference from the trivial upper bound of the optimal value.
2 Preliminaries
2.1 Definitions
We refer the readers to [25] for basic concepts and notation on graphs. In this paper, a graph means a directed graph or an undirected graph. We call a directed graph simply a digraph.
For a graph $G$, we denote by $V(G)$ and $E(G)$ the vertex set and the edge set of $G$, respectively. A subgraph of $G$ is a graph $G^{\prime}$ such that $V(G^{\prime})\subseteq V(G)$, $E(G^{\prime})\subseteq E(G)$, and each edge $e\in E(G^{\prime})$ is between two vertices in $V(G^{\prime})$. A subgraph $G^{\prime}$ of $G$ is induced by a vertex set $X\subseteq V(G)$, denoted by $G[X]$, if $V(G^{\prime})=X$ and $E(G^{\prime})$ is the set of edges in $E(G)$ whose end vertices are both in $X$. We also define $G-X\coloneqq G[V\setminus X]$ for $X\subseteq V$. A subgraph $G^{\prime}$ of $G$ is spanning if $V(G^{\prime})=V(G)$. We often do not distinguish between a spanning subgraph $G^{\prime}$ and its edge set $E(G^{\prime})$.
A stable set in an undirected graph $G$ is a vertex set $X\subseteq V(G)$ such that no edge in $E(G)$ has its both end vertices in $X$. An undirected graph $G=(V,E)$ is bipartite if there exists a bipartition $(V^{+},V^{-})$ of $V$ such that $V^{+}\cap V^{-}=\emptyset$, $V^{+}\cup V^{-}=V$, and both $V^{+}$ and $V^{-}$ are stable sets in $G$ (possibly $V^{+}$ or $V^{-}$ is empty). To specify the bipartition, we denote by $G=(V^{+},V^{-};E)$. For a vertex set $X$ on one side (i.e., $X\subseteq V^{+}$ or $X\subseteq V^{-}$), let $\Gamma_{G}(X)$ denote the neighborhood of $X$, i.e., the set of vertices adjacent to some vertex in $X$. A bipartite graph $G=(V^{+},V^{-};E)$ is balanced if $|V^{+}|=|V^{-}|$.
For a digraph $G=(V,E)$, the underlying graph is an undirected graph $\overline{G}=(V,\overline{E})$ defined as follows. For each directed edge $e=(u,v)\in E$, we denote by $\overline{e}=\{u,v\}$ the corresponding undirected edge. We then define $\overline{F}\coloneqq\{\overline{e}\mid e\in F\}$ for each $F\subseteq E$.
For each edge $e=\{u,v\}$ in a bipartite graph $G=(V^{+},V^{-};E)$ such that $u\in V^{+}$ and $v\in V^{-}$, we denote its orientations by $\overrightarrow{e}=(u,v)$ and $\overleftarrow{e}=(v,u)$. Similarly, for each $F\subseteq E$, we define $\overrightarrow{F}$ and $\overleftarrow{F}$ as the sets of the orientations $\overrightarrow{e}$ and $\overleftarrow{e}$, respectively, of the edges $e\in F$.
An undirected graph is said to be connected if for each pair of vertices, there exists a path between them. A digraph is said to be strongly connected if the same condition (where a path means a directed path) holds, and just connected if the underlying graph is connected.
For an undirected graph $G$ and a vertex $v\in V(G)$, we denote by $\deg_{G}(v)$ the degree of $v$, which is the number of edges incident to $v$. Similarly, for a digraph $G$ and a vertex $v\in V(G)$, we denote by $\deg_{G}^{\mathrm{in}}(v)$ and $\deg_{G}^{\mathrm{out}}(v)$ the in-degree and out-degree of $v$, respectively, which are the numbers of edges entering and leaving $v$. An in-arborescence (or out-arborescence) is a connected digraph in which exactly one vertex, called the root, has out-degree (or in-degree, respectively) 00 and all other vertices have out-degree (or in-degree, respectively) $1$. An in- or out-arborescence having its root $r$ is said to be rooted at $r$, and is also called an $r$-in-arborescence or an $r$-out-arborescence, respectively.
A matching in an undirected graph $G$ is an edge subset in which no two edges share a vertex. If a matching covers all the vertices in $G$, it is called a perfect matching. For a perfect matching $M$ in a balanced bipartite graph $G=(V^{+},V^{-};E)$, the auxiliary graph of $G$ with respect to $M$ is a digraph $G_{M}\coloneqq(V;\overrightarrow{E}\cup\overleftarrow{M})$.
The DM-decomposition [7, 6] of a bipartite graph is formally defined by the distributive lattice formed by the minimizers of a submodular function, and a bipartite graph is called DM-irreducible if its DM-decomposition consists of a single component. We here employ its characterization as an alternative “definition” (cf. [3, Section 2.2]). [^1] The equivalent condition is also called elementary, e.g., in [23, 11].
Lemma 2.1 (cf. [11, Theorem 2.4.18]).
For a balanced bipartite graph $G$, the following statements are equivalent.
  1. 1.
    $G$ is DM-irreducible.
  2. 2.
    $G$ has a perfect matching $M$ such that $G_{M}$ is strongly connected.
  3. 3.
    $G$ has a perfect matching, and for any perfect matching $M$ in $G$, $G_{M}$ is strongly connected.
Suppose that a graph $G$ is associated with edge weight $w\colon E\to\mathbb{R}$. We define the weight of each edge set $F\subseteq E$ as $w(F)\coloneq\sum_{e\in F}w(e)$.
2.2 Problems and Our Results
We formally state the problems dealt with in this paper. Recall that DMISS for the unbalanced bipartite graphs is tractable [3], and it is not considered in this paper.
Problem (Strongly Connected Spanning Subgraph (SCSS)).
Input:
A strongly connected digraph $G=(V,E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
Goal:
Find a minimum-weight strongly connected spanning subgraph of $G$.
Problem (DM-Irreducible Spanning Subgraph (DMISS)).
Input:
A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
Goal:
Find a minimum-weight DM-irreducible spanning subgraph of $G$.
It is easy to observe that any inclusion-wise minimal strongly connected spanning subgraph of a digraph on $n$ vertices consists of at most $22n-$ 2 edges (by Lemma 3.1). Similarly, it is easy to observe that any inclusion-wise minimal DM-irreducible spanning subgraph of a balanced bipartite graph on $2n$ vertices ($n$ vertices on each side) consists of at most $23n-$ 2 edges (by Lemma 3.2). The following problems are the unweighted version of the above two problems parameterized by the difference from these trivial upper bounds (note that it is obvious from the definition that $|V|$ edges are necessary for either problem).
Problem (UnweightedSCSS).
Input:
A strongly connected digraph $G=(V,E)$.
Parameter:
An integer $k\leq|V|-2$.
Goal:
Test whether $G$ has a strongly connected spanning subgraph with at most $2|V|-2-k$ edges or not.
Problem (UnweightedDMISS).
Input:
A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$.
Parameter:
An integer $k\leq|V^{+}|-2$.
Goal:
Test whether $G$ has a DM-irreducible spanning subgraph with at most $3|V^{+}|-2-k$ edges or not.
The results shown in this paper are summarized as follows, where $n=|V^{+}|=|V^{-}|$.
Theorem 2.2.
There exists a $2$-approximation algorithm for DMISS that runs in $O(n^{3})$ time.
Theorem 2.3.
There exists an FPT algorithm for UnweightedDMISS.
3 Approximation Algorithms for DMISS
In this section, we prove Theorem 2.2 by presenting a $2$-approximation algorithm for DMISS that runs in $O(n^{3})$ time.
In Section 3.1, as a base of our result, we describe a $2$-approximation algorithm for SCSS by Frederickson and Jájá [14]. In Section 3.2, we show a $3$-approximation algorithm for DMISS obtained by a naive extension of the result of Frederickson and Jájá. In Section 3.3, we improve it to a $2$-approximation algorithm with the aid of matroid intersection.
3.1 2-approximation Algorithm for SCSS
The algorithm is based on the following characterization of strongly connected digraphs.
Lemma 3.1.
For a digraph $G$, the following statements are equivalent.
  1. 1.
    $G$ is strongly connected.
  2. 2.
    For some vertex $r$, $G$ has a spanning in-arborescence and out-arborescence rooted at $r$.
  3. 3.
    For any vertex $r$, $G$ has a spanning in-arborescence and out-arborescence rooted at $r$.
For the input $(G=(V,E),w)$ of SCSS, the algorithm arbitrarily chooses a vertex $r\in V$ and finds a minimum-weight spanning in-arborescence $A_{\mathrm{in}}$ and out-arborescence $A_{\mathrm{out}}$ rooted at $r$; the existence of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ follows from Lemma 3.1 $(1\Rightarrow 3)$. This can be done in $O(m+n\log n)$ time [15], where $n=|V|$ and $m=|E|$. By Lemma 3.1 $(2\Rightarrow 1)$, the union of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ results in a strongly connected spanning subgraph of $G$, and the algorithm outputs this. Since any optimal solution has a spanning in-arborescence and out-arborescence rooted at $r$ by Lemma 3.1 $(1\Rightarrow 3)$ and their weights are at least those of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$, respectively, the weight of the output solution is at most twice the optimal value.
3.2 3-approximation Algorithm for DMISS
A naive extension of the above result leads to a $3$-approximation algorithm for DMISS. First, Lemma 3.1 is straightforwardly extended as follows (recall Lemma 2.1).
Lemma 3.2.
For a balanced bipartite graph $G$, the following statements are equivalent.
  1. 1.
    $G$ is DM-irreducible.
  2. 2.
    $G$ has a perfect matching $M$ and a vertex $r$ such that $G_{M}$ has a spanning in-arborescence and out-arborescence rooted at $r$.
  3. 3.
    $G$ has a perfect matching, and for any perfect matching $M$ and any vertex $r$, $G_{M}$ has a spanning in-arborescence and out-arborescence rooted at $r$.
To achieve $3$-approximation for DMISS, we simply divide the problem into two tasks: finding a minimum-weight perfect matching $M$ in $(G,w)$ and solving SCSS with the input $(G_{M},w_{M})$, where $w_{M}\colon E(G_{M})\to\mathbb{R}_{\geq 0}$ is defined as
$w_{M}(e)\coloneqq\begin{cases}w(\overline{e})&(\overline{e}\not\in M),\\ 0&(\overline{e}\in M).\end{cases}$ (1)
For the first task, polynomial-time exact algorithms initiated by the Hungarian method [22] are well-known. For the second task, we apply the $2$-approximation algorithm described in the previous section. We then output the union of the solutions (for the second solution, we take the set of corresponding edges in $G$). This intuitively leads to a $3$-approximation algorithm for DMISS (as $1+2=3$), which can be formally shown as follows.
Lemma 3.3.
There exists a polynomial-time $3$-approximation algorithm for DMISS.
Proof.
Let $M$ be a minimum-weight perfect matching in $(G,w)$ (for the first task), and $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ be a minimum-weight spanning in-arborescence and out-arborescence in $(G_{M},w_{M})$ (for the second task), respectively. Also, let $S_{\mathrm{opt}}$ be an optimal solution and $M_{\mathrm{opt}}$ be a perfect matching in $S_{\mathrm{opt}}$. We then have $w(M)\leq w(M_{\mathrm{opt}})\leq w(S_{\mathrm{opt}})$.
Consider a subgraph $G^{\prime}=(V^{+},V^{-};S_{\mathrm{opt}}\cup M)$ of $G$, which is DM-irreducible as so is $S_{\mathrm{opt}}$. By Lemma 3.2, the auxiliary graph $G^{\prime}_{M}$ has a spanning in-arborescence and out-arborescence rooted at any vertex. Let $r$ be the root of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$, and $A_{\mathrm{in}}^{\prime}$ and $A_{\mathrm{out}}^{\prime}$ be any spanning in-arborescence and out-arborescence of $G^{\prime}_{M}$ rooted at $r$, respectively. We then have
$w((\overline{A_{\mathrm{in}}}\cup\overline{A_{\mathrm{out}}})\setminus M)$ $=w_{M}(A_{\mathrm{in}}\cup A_{\mathrm{out}})$ (definition of $w_{M}$)
$\leq w_{M}(A_{\mathrm{in}})+w_{M}(A_{\mathrm{out}})$ ($w_{M}$ is nonnegative)
$\leq w_{M}(A_{\mathrm{in}}^{\prime})+w_{M}(A_{\mathrm{out}}^{\prime})$ ($G^{\prime}_{M}$ is a subgraph of $G_{M}$)
$=w(\overline{A_{\mathrm{in}}^{\prime}}\setminus M)+w(\overline{A_{\mathrm{out}}^{\prime}}\setminus M)$ (definition of $w_{M}$)
$\leq 2w(S_{\mathrm{opt}})$ $(E(G^{\prime})\setminus M\subseteq S_{\mathrm{opt}}).$
Thus, the weight of the output solution is
$w(M\cup\overline{A_{\mathrm{in}}}\cup\overline{A_{\mathrm{out}}})=w(M)+w((\overline{A_{\mathrm{in}}}\cup\overline{A_{\mathrm{out}}})\setminus M)\leq 3w(S_{\mathrm{opt}}).\qed$
3.3 2-approximation Algorithm for DMISS
We improve the above $3$-approximation algorithm with the aid of matroid intersection.
We first give an easy but important observation. In short, for any $r\in V^{+}$, all the matching edges are included in every spanning $r$-in-arborescence in the auxiliary graph.
Lemma 3.4.
Let $G=(V^{+},V^{-};E)$ be a DM-irreducible balanced bipartite graph, $M$ be a perfect matching in $G$, and $r\in V^{+}$. Then, any spanning $r$-in-arborescence of $G_{M}$ contains all the edges in $\overleftarrow{M}$.
Proof.
Let $A_{\mathrm{in}}$ be a spanning $r$-in-arborescence of $G_{M}$. Then, for any vertex $v\in V^{-}$, we have $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)=1$. In $G_{M}=(V^{+},V^{-};\overrightarrow{E}\cup\overleftarrow{M})$, exactly one edge leaves each vertex in $V^{-}$, which is in $\overleftarrow{M}$. Thus, $A_{\mathrm{in}}$ must contain all the edges in $\overleftarrow{M}$. ∎
In the previous algorithm, we essentially solve three problems: finding a minimum-weight perfect matching, a minimum-weight spanning in-arborescence, and a minimum-weight spanning out-arborescence. Based on Lemma 3.4, we merge the first two into a single task as follows: finding a minimum-weight spanning tree $T$ in $(G,w)$ among those having a perfect matching $M$ such that $\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is a spanning in-arborescence of $G_{M}$.
The main observation is the following characterization of such spanning trees.
Lemma 3.5.
For a spanning tree $T$ of a balanced bipartite graph $G$, the following statements are equivalent.
  1. 1.
    $T$ has a perfect matching $M$ in $G$ and $\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is a spanning in-arborescence of $G_{M}$.
  2. 2.
    There exists a vertex $r\in V^{+}$ such that $\deg_{T}(r)=1$ and $\deg_{T}(v)=2$ for every $v\in V^{+}\setminus\{r\}$.
Proof.
$(1\Rightarrow 2)$  Let $A_{\mathrm{in}}=\overrightarrow{T\setminus M}\cup\overleftarrow{M}$. Since $M$ is a perfect matching in $G$, $\deg_{A_{\mathrm{in}}}^{\mathrm{in}}(v)=1$ for every $v\in V^{+}$ and $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)=1$ for every $v\in V^{-}$. Also, since $A_{\mathrm{in}}$ is a spanning in-arborescence, $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)=1$ for every $v\in V^{+}\setminus\{r\}$ and $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(r)=0$, where $r$ is the root of $A_{\mathrm{in}}$. As $\deg_{T}(v)=\deg_{A_{\mathrm{in}}}^{\mathrm{in}}(v)+\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)$ for every vertex $v$, we are done.
$(2\Rightarrow 1)$  We first show that $T$ has a perfect matching. By Hall’s theorem [16], it suffices to show that for any vertex set $X^{+}\subseteq V^{+}$, we have $|X^{+}|\leq|\Gamma_{T}(X^{+})|$. Suppose to the contrary that there exists a vertex set $X^{+}\subseteq V^{+}$ such that $|X^{+}|>|\Gamma_{T}(X^{+})|$. Then, the subgraph formed by the edges in $T$ intersecting $X^{+}$ has at most $12|X^{+}|-$ 1 vertices and at least $12|X^{+}|-$ 1 edges (due to the degree condition), so it must contain a cycle, a contradiction.
Let $M$ be the perfect matching in $T$ (note that it is unique). We next show that $A_{\mathrm{in}}=\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is an in-arborescence. As $T$ is a tree, it suffices to show that $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)=1$ for every vertex $v\in V\setminus\{r\}$, where $r\in V^{+}$ is the vertex specified in Statement 2. By definition, every vertex $v\in V^{-}$ has a unique outgoing edge, which is in $\overleftarrow{M}$, and hence $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)=1$. Also, every vertex $v\in V^{+}$ has a unique incoming edge, which is in $\overleftarrow{M}$, and hence $\deg_{A_{\mathrm{in}}}^{\mathrm{in}}(v)=1$. Thus, for every vertex $v\in V^{+}\setminus\{r\}$, we have $\deg_{A_{\mathrm{in}}}^{\mathrm{out}}(v)=\deg_{T}(v)-\deg_{A_{\mathrm{in}}}^{\mathrm{in}}(v)=2-1=1$, and we are done. ∎
Let us call a spanning tree satisfying Statement 2 in Lemma 3.5 strongly balanced, and say that the vertex $r$ is the root and the tree is rooted at $r$. By Lemma 3.2, every DM-irreducible graph has a strongly balanced spanning tree, and then we formulate the first task as follows.
Problem (Strongly Balanced Spanning Tree (SBST)).
Input:
A DM-irreducible bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
Goal:
Find a minimum-weight strongly balanced spanning tree of $G$.
For the sake of simplicity of the following discussion, we consider the following problem, which is polynomially equivalent to the above problem and actually enough for our purpose.
Problem (Strongly Balanced Spanning Tree with Root (SBSTR)).
Input:
A DM-irreducible bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$, and $r\in V^{+}$.
Goal:
Find a minimum-weight strongly balanced spanning tree of $G$ rooted at $r$.
Our proposed algorithm for DMISS is described as follows.
Step 1.
Choose an arbitrary vertex $r\in V^{+}$ as a root, and solve SBSTR with the input $(G,w,r)$. Let $T$ be the obtained solution, and $M$ be the perfect matching in $T$.
Step 2.
Find a minimum-weight spanning $r$-out-arborescence $A_{\mathrm{out}}$ in $(G_{M},w_{M})$ (recall that $w_{M}$ is defined as (1)).
Step 3.
Output $T\cup\overline{A_{\mathrm{out}}}$.
The output is indeed a DM-irreducible spanning subgraph (by Lemma 3.2). To complete the proof of Theorem 2.2, we analyze the computational time and the approximation ratio.
Computational Time
As we already mentioned, the second task (finding a minimum-weight spanning arborescence) can be done in $O(m+n\log n)$ time [15]. The following observation (which was also observed in [13]) leads to an $O(n^{3})$-time algorithm for the first task (solving SBSTR), which completes the analysis. We omit the basics on matroids, for which we refer the readers to [25].
Lemma 3.6 (cf. [13, Lemma 5]).
SBST and SBSTR are formulated as the weighted matroid intersection problem, i.e., the set of feasible solutions can be represented as the set of common bases of two matroids.
Proof.
We first introduce two matroids. Let $\mathbf{M}_{1}=(E,\mathcal{I}_{1})$ be the cycle matroid of $G$, i.e.,
[…]
and $\mathbf{M}_{2}=(E,\mathcal{I}_{2})$ be a partition matroid defined by
$\mathcal{I}_{2}=\{F\subseteq E\mid\deg_{F}(v)\leq 2\ (\forall v\in V^{+})\}.$
Note that $V^{+}$ is a stable set in $G$ by definition.
In SBST, the set of feasible solutions is represented as the set of common bases of $\mathbf{M}_{1}$ and the truncation of $\mathbf{M}_{2}$ by $12|V^{+}|-$ 1, whose independent set family is $\{F\in\mathcal{I}_{2}\mid|F|\leq 2|V^{+}|-1\}$. In SBSTR, the set of feasible solutions is represented as the set of common bases of $\mathbf{M}_{1}$ and a slightly modified partition matroid $\mathbf{M}^{\prime}_{2}=(E,\mathcal{I}^{\prime}_{2})$ such that
$\mathcal{I}^{\prime}_{2}=\{F\subseteq E\mid\deg_{F}(r)\leq 1\text{ and }\deg_{F}(v)\leq 2\ (\forall v\in V^{+}\setminus\{r\})\}.\qed$
Lemma 3.6 implies that SBST and SBSTR can be solved in polynomial time [^2] In contrast, it was recently shown that finding a strongly balanced spanning tree in a nonbipartite graph is NP-hard [4]. with the aid of matroid intersection algorithms. In particular, Brezovec, Cornuéjols, and Glover [5] proposed an $O(\tilde{n}\tilde{m}+\tilde{k}\tilde{n}^{2})$-time algorithm for the following special case: one matroid is a cycle matroid of a graph and the other is a partition matroid whose partition classes are induced by a stable set of the graph (i.e., each partition class is the set of edges incident to a vertex in the stable set), where $\tilde{n}$ and $\tilde{m}$ are the numbers of vertices and of edges in the graph, respectively, and $k$ is the size of the stable set inducing the partition classes. By the proof of Lemma 3.6, SBSTR completely fits this situation, and hence we can exactly solve it in $O(n^{3})$ time, where we have $\tilde{k}=|V^{+}|=n=\frac{\tilde{n}}{2}$ and $\tilde{m}\leq n^{2}$.
Remark 3.7.
It seems not easy to essentially improve the computational time. Both SBST and SBSTR actually include two fundamental special cases of the weighted matroid intersection problem: finding a minimum-weight perfect matching in a bipartite graph, and finding a minimum-weight spanning arborescence in a digraph. In particular, for the former problem, one of the best-known bounds in general is $O(n(m+n\log n))$ [8], which is not better than $O(n^{3})$ when the graph is dense (i.e., $m=\Theta(n^{2})$).
We sketch that SBSTR indeed includes the two problems. For a balanced bipartite graph $G=(V^{+},V^{-};E)$, add two vertices $s^{+},s^{-}$ and $|V^{+}|+1$ edges $\{v,s^{-}\}$ $(v\in V^{+}\cup\{s^{+}\})$ of weight 00, and add a sufficiently large value (uniformly) to the weights of all the edges in $E$. Then, a minimum-weight perfect matching in $G$ corresponds to a minimum-weight strongly balanced spanning tree of the resulting graph rooted at $s^{+}$, which must use all the additional edges of weight 00.
Also, for a digraph $G=(V,E)$, split each vertex $v\in V$ into two vertices $v^{+}$ and $v^{-}$ with an edge $\{v^{+},v^{-}\}$ of weight 00, replace each edge $(u,v)$ with $\{u^{+},v^{-}\}$ of the same weight plus a sufficiently large value (the same for all the edges). Then, the resulting graph $G^{\prime}$ is indeed a balanced bipartite graph, and a minimum-weight spanning in-arborescence of $G$ rooted at $r\in V$ corresponds to a minimum-weight strongly balanced spanning tree of $G^{\prime}$ rooted at $r^{+}$, which must use all the edges $\{v^{+},v^{-}\}$ of weight 00.
Approximation Ratio
Let $S_{\mathrm{opt}}$ be an optimal solution. We show $w(T\cup\overline{A_{\mathrm{out}}})=w(T)+w_{M}(A_{\mathrm{out}})\leq 2w(S_{\mathrm{opt}})$ (recall that $w_{M}\colon E(G_{M})\to\mathbb{R}_{\geq 0}$ is defined by (1)). The following two claims imply this.
Claim 3.8.
$w(T)\leq w(S_{\mathrm{opt}})$.
Proof.
By Lemmas 3.2 and 3.4, $S_{\mathrm{opt}}$ has a perfect matching $M_{\mathrm{opt}}$, and there exists a spanning $r$-in-arborescence $A_{\mathrm{in}}^{\prime}$ of $G_{M_{\mathrm{opt}}}$ such that $\overleftarrow{M_{\mathrm{opt}}}\subseteq A_{\mathrm{in}}^{\prime}\subseteq\overrightarrow{S_{\mathrm{opt}}}\cup\overleftarrow{M_{\mathrm{opt}}}$. Since $T$ is optimal and $\overline{A_{\mathrm{in}}^{\prime}}$ is feasible in SBSTR with the input $(G,w,r)$, we have $w(T)\leq w(\overline{A_{\mathrm{in}}^{\prime}})\leq w(S_{\mathrm{opt}})$. ∎
Claim 3.9.
$w_{M}(A_{\mathrm{out}})\leq w(S_{\mathrm{opt}})$.
Proof.
As with the proof of Lemma 3.3, consider a subgraph $G^{\prime}=(V^{+},V^{-};S_{\mathrm{opt}}\cup M)$ of $G$. Then, by Lemma 3.2, $G^{\prime}_{M}$ has a spanning $r$-out-arborescence $A_{\mathrm{out}}^{\prime}$, and hence
$w_{M}(A_{\mathrm{out}})\leq w_{M}(A_{\mathrm{out}}^{\prime})=w(\overline{A_{\mathrm{out}}^{\prime}}\setminus M)\leq w(S_{\mathrm{opt}}).\qed$
Figure 1: A tight example of our $2$-approximation algorithm for DMISS.
Remark 3.10.
The above analysis is tight. Consider the input $(G,w)$ shown in Figure 1, which consists of
  • $2(2\ell+3)$ vertices $v_{i}^{+},v_{i}^{-}$ $(i=0,1,\dots,2\ell+2)$,
  • $32\ell+$ 3 edges $\{v_{i}^{+},v_{i}^{-}\}$ $(i=0,1,\dots,2\ell+2)$ of weight 00,
  • $\ell$ edges $\{v_{2i-1}^{+},v_{2i}^{-}\}$ $(i=1,2,\dots,\ell)$ of weight $1+\varepsilon$,
  • $\ell$ edges $\{v_{2i}^{+},v_{2i+1}^{-}\}$ $(i=1,2,\dots,\ell)$ of weight $\varepsilon$,
  • $\ell$ edges $\{v_{2i-1}^{+},v_{2i+1}^{-}\}$ $(i=1,2,\dots,\ell)$ of weight $1$,
  • two edges $\{v_{0}^{+},v_{1}^{-}\}$ and $\{v_{2\ell+1}^{+},v_{2\ell+2}^{-}\}$ of weight $\varepsilon$, and
  • an edge $\{v_{2\ell+2}^{+},v_{0}^{-}\}$ of weight $1+\varepsilon$,
where $\ell$ is a positive integer and $\varepsilon\coloneqq\frac{1}{\ell}>0$.
While a Hamiltonian cycle $(v_{0}^{-},v_{0}^{+},v_{1}^{-},v_{1}^{+},\dots,v_{2\ell+2}^{-},v_{2\ell+2}^{+},v_{0}^{-})$ illustrated as in (b) is a feasible solution, the output of our algorithm is the whole graph as follows. First, the edges of weight at most $1$ form a spanning tree of $G$, which is of minimum weight (consider Kruskal’s algorithm [21]). This minimum-weight spanning tree is strongly balanced with root $r=v_{2\ell+2}^{+}$, and hence it is eligible for $T$ in Step 1. The edges of weight 00 form a perfect matching $M$, and any spanning $r$-out-arborescence of $G_{M}$ (any candidate of $A_{\mathrm{out}}$ in Step 2) must contain all the remaining edges, which are of weight $1+\varepsilon$ (as they are the only edges entering $v_{2i}$ $(i=1,2,\dots,\ell)$ in $G_{M}$).
Thus, the approximation ratio of our algorithm is at least
$\frac{w(T\cup\overline{A_{\mathrm{out}}})}{w(S_{\mathrm{opt}})}\geq\frac{\ell(2+2\varepsilon)+2\varepsilon+1+\varepsilon}{\ell(1+2\varepsilon)+2\varepsilon+1+\varepsilon}=\frac{2\ell^{2}+3\ell+3}{\ell^{2}+3\ell+3}\to 2\quad(\ell\to\infty).$
4 FPT Algorithm for Unweighted DMISS
In this section, we prove Theorem 2.3 by presenting an FPT algorithm for UnweightedDMISS.
In Section 4.1, as a base of our result, we sketch the idea of an FPT algorithm for UnweightedSCSS proposed by Bang-Jensen and Yeo [2]. In Section 4.2, by extending it, we propose an FPT algorithm for UnweightedDMISS.
4.1 FPT Algorithm for Unweighted SCSS
The algorithm utilizes another well-known characterization of strongly connected digraphs. For a digraph $G=(V,E)$, an ear decomposition of $G$ is a sequence $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of edge-disjoint paths and cycles in $G$ with $E=\bigcup_{i=0}^{f}E(P_{i})$ such that
  • $P_{0}$ is a cycle, and
  • $P_{i}$ is either a path between two different vertices in $\bigcup_{j=0}^{i-1}V(P_{j})$ whose inner vertices are in $V\setminus\bigcup_{j=0}^{i-1}V(P_{j})$, or a cycle intersecting exactly one vertex in $\bigcup_{j=0}^{i-1}V(P_{j})$.
Each path or cycle $P_{i}$ is called an ear. An ear is called long if it is of length at least $3$.
Lemma 4.1 (cf. [11, Theorem 2.2.1]).
A digraph is strongly connected if and only if it has an ear decomposition.
The idea of the FPT algorithm for UnweightedSCSS in [2] is sketched as follows. It first computes an ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of the input graph $G$ so that the set of long ears is inclusion-wise maximal (i.e., so that no more long ear can be taken in addition to them, but the number of long ears in $\mathcal{E}$ is not necessarily maximized). If the number of long ears is large enough (depending on the parameter $k$), then the ear decomposition $\mathcal{E}$ gives a strongly connected spanning subgraph that is small enough by removing all the ears of length $1$, and the answer is yes. Otherwise, let $X$ be the set of vertices intersecting a long ear in $\mathcal{E}$. Under the assumption that $\mathcal{E}$ does not immediately give a solution, we can show that $|X|=O(k)$. Also, by the maximality of long ears in $\mathcal{E}$, the vertices in $V\setminus X$ are added by ears of length $2$, and this is basically true for some minimum strongly connected spanning subgraph. Thus, we can test the existence of a solution by an exhaustive search on the subgraphs of $G$ on the vertex set $X$ (plus a small subset), where we have only $2^{O(k^{2})}$ candidates.
4.2 FPT Algorithm for Unweighted DMISS
First, we give an alternative characterization of DM-irreducible balanced bipartite graphs by ear decompositions. An ear decomposition of a bipartite graph $G=(V^{+},V^{-};E)$ is defined in the same way as above, and we call it odd proper if
  • $P_{0}$ is a cycle of length at least $4$, and
  • each $P_{i}$ $(i=1,2,\dots,f)$ is a path of odd length, which implies that one end vertex of $P_{i}$ is in $V^{+}$ and the other is in $V^{-}$.
Lemma 4.2 (cf. [11, Theorem 2.4.17] or [23, Theorem 4.1.6]).
A bipartite graph on at least four vertices is balanced and DM-irreducible if and only if it has an odd proper ear decomposition.
Proof.
We give a proof to make this paper self-contained.
Suppose that a bipartite graph $G=(V^{+},V^{-};E)$ with $|V|\geq 4$ is balanced and DM-irreducible. By Lemma 2.1, there exists a perfect matching $M$ in $G$ and $G_{M}$ is strongly connected. Since $|V|\geq 4$ and $G_{M}$ is connected, there exists an edge $e_{0}=\{u_{0},v_{0}\}\in E\setminus M$, where $u_{0}\in V^{+}$ and $v_{0}\in V^{-}$. As $G_{M}$ is strongly connected, it contains a path from $v_{0}$ to $u_{0}$, which forms a cycle of length at least $4$ together with the edge $\overrightarrow{e_{0}}=(u_{0},v_{0})$. Let $(P_{0},P_{1},\dots,P_{f})$ be an arbitrary ear decomposition of $G_{M}$ such that $P_{0}$ is this cycle (the existence follows from Lemma 4.1 with shrinking $P_{0}$ into a single vertex). We then see by induction on $i=0,1,\dots,f-1$ that
  • for every $e=\{u,v\}\in M$, either $\{u,v\}\cap\bigcup_{j=0}^{i}V(P_{j})=\emptyset$ or $\{u,v\}\subseteq\bigcup_{j=0}^{i}V(P_{j})$ and $\overleftarrow{e}=(v,u)\in\bigcup_{j=0}^{i}E(P_{j})$, and
  • $P_{i+1}$ starts and ends with edges in $\overrightarrow{E}$, which implies that it is of odd length.
Thus, we are done by taking the corresponding ear decomposition of the underlying graph (with removing each ear consisting of a single edge $\overrightarrow{e}\in\overrightarrow{M}$, which is duplicated).
Suppose that a bipartite graph $G$ with $|V(G)|\geq 4$ has an odd proper ear decomposition $(P_{0},P_{1},\dots,P_{f})$. We then see by induction on $i=0,1,\dots,f$ that the subgraph $G_{i}$ of $G$ defined by $V(G_{i})\coloneqq\bigcup_{j=0}^{i}V(P_{j})$ and $E(G_{i})\coloneqq\bigcup_{j=0}^{i}E(P_{j})$ is balanced and DM-irreducible (it is straightforward by Lemma 2.1), which concludes that $G$ itself is balanced and DM-irreducible. ∎
An ear is called trivial if it consists of a single edge, and nontrivial otherwise. Also, an ear is called long if it is of length at least $5$. [^3] This redefinition is reasonable by considering the relation between the strong connectivity of a digraph and the DM-irreducibility of a balanced bipartite graph (cf. [3, Section 2.3])
We start to describe our FPT algorithm for UnweightedDMISS. Let $G=(V^{+},V^{-};E)$ be the input DM-irreducible balanced bipartite graph, where $n=|V^{+}|\geq 2$ and $m=|E|$. We first find an odd proper ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ such that
  • $\{P_{1},P_{2},\dots,P_{s}\}$ is the set of long ears in $\mathcal{E}$, which is inclusion-wise maximal (i.e., no more long ear can be taken in addition to them, where $s$ is not necessarily maximized), and
  • $\{P_{1},P_{2},\dots,P_{r}\}$ is the set of nontrivial ears in $\mathcal{E}$ (note that $s\leq r\leq f$).
This can be done in polynomial time. [^4] For example, naively, we can start with an arbitrary cycle $P_{0}$ of length at least $4$, and greedily add a long ear as far as possible; a long ear can be found by an exhaustive search of the first and last three vertices, where the number of candidates is $O(n^{6})$.
For each $i=0,1,\dots,f$, we define a subgraph $G_{i}$ of $G$ by $V(G_{i})\coloneqq\bigcup_{j=0}^{i}V(P_{j})$ and $E(G_{i})\coloneqq\bigcup_{j=0}^{i}E(P_{j})$. Let $X=V(G_{s})$ and $Y=V\setminus X$. We then observe the following two claims.
Claim 4.3.
$G_{r}$ is a DM-irreducible spanning subgraph of $G$, which has $(3n-2)-\left(\frac{|X|}{2}-s-2\right)$ edges.
Proof.
Since $V(G_{r})=V(G)$, by Lemma 4.2, $G_{r}$ is a DM-irreducible spanning subgraph of $G$. By definition, we have
$|V(G_{r})|$ $=|V(P_{0})|+\sum_{i=1}^{r}(|V(P_{i})|-2)=|V(P_{0})|+\sum_{i=1}^{r}(|E(P_{i})|-1),$
$|E(G_{r})|$ $=|E(P_{0})|+\sum_{i=1}^{r}|E(P_{i})|=|V(P_{0})|+\sum_{i=1}^{r}|E(P_{i})|,$
and hence $|E(G_{r})|=|V(G_{r})|+r=2n+r=(3n-2)-(n-r-2)$. We also have
$|X|=2n-|Y|=2n-\sum_{i=s+1}^{r}(|E(P_{i})|-1)=2n-2(r-s),$
and we are done. ∎
By Claim 4.3, if $k\leq\frac{|X|}{2}-s-2$, then $(G,k)$ is a yes-instance. We also have $|X|\geq 4s+4$ by definition of long ears. In what follows, we assume $k\geq\frac{|X|}{2}-s-1\geq\frac{|X|}{4}$, i.e., $|X|\leq 4k$.
Claim 4.4.
$E(G[Y])$ is a perfect matching.
Proof.
By definition, $P_{s+1},P_{s+2},\dots,P_{r}$ are paths of length $3$ and the two inner vertices of each path are in $Y$. Thus, $E(G[Y])$ indeed has a perfect matching
$\{\{y_{i}^{+},y_{i}^{-}\}\mid\text{$\{y_{i}^{+},y_{i}^{-}\}$ is the middle edge of the ear $P_{i}$ $(s+1\leq i\leq r)$}\}.$
By the maximality of long ears, we can see that $E(G[Y])$ contains no other edge as follows.
Suppose to the contrary that $E(G[Y])$ contains another edge $e=\{y_{i}^{+},y_{j}^{-}\}$ with $i\neq j$. We pick such an edge minimizing $i+j$, and assume $i<j$ by symmetry. Then, $G_{j}$ has a path $Q_{i}$ of length $2$ between $y_{i}^{+}$ and some vertex $x^{+}\in X\cap V^{+}$ (a part of the ear $P_{i}$) and a path $Q_{j}$ of length at least $2$ between $y_{j}^{-}$ and some vertex $x^{-}\in X\cap V^{-}$ starting with the edge $\{y_{j}^{+},y_{j}^{-}\}$ whose inner vertices are in $Y$. If $Q_{i}$ and $Q_{j}$ are disjoint, then $Q_{i}$, $Q_{j}$, and $e$ form a long ear that can be added to $G_{s}$ (just after $P_{s}$ in $\mathcal{E}$), which contradicts the maximality of long ears. Otherwise, $Q_{j}$ start with the edges $\{y_{j}^{+},y_{j}^{-}\}$ and $\{y_{j}^{+},y_{i}^{-}\}$ (otherwise, there exists another edge $e^{\prime}=\{y_{i^{\prime}}^{+},y_{j^{\prime}}^{-}\}$ with $i^{\prime}\neq j^{\prime}$ and $i^{\prime}+j^{\prime}<i+j$, a contradiction). In this case, the first two edges of $Q_{j}$, $e$, and the two end edges in the ear $P_{i}$ form a long ear that can be added to $G_{s}$ (just after $P_{s}$ in $\mathcal{E}$), which again contradicts the maximality of long ears. ∎
We consider a bipartite graph $\tilde{H}$ with bipartition $(\tilde{Z},\tilde{Y})$ defined as follows:
$\tilde{Y}$$\coloneqq\{\tilde{y}\coloneqq\{y^{+},y^{-}\}\mid y^{+},y^{-}\in Y,\ \{y^{+},y^{-}\}\in E\},$
$\tilde{Z}$$\coloneqq\{\tilde{z}\coloneqq\{z^{+},z^{-}\}\mid\exists\tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y},\ \{\{z^{+},y^{-}\},\{z^{-},y^{+}\}\}\subseteq E\},$
$E(\tilde{H})$$\coloneqq\{\{\tilde{z},\tilde{y}\}\mid\tilde{z}=\{z^{+},z^{-}\}\in\tilde{Z},\ \tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y},\ \{\{z^{+},y^{-}\},\{z^{-},y^{+}\}\}\subseteq E\}.$
Note that each vertex in $\tilde{H}$ is a pair of vertices in $G$, one of which is in $V^{+}$ and the other is in $V^{-}$.
Let $\tilde{Z}_{\mathrm{out}}\subseteq\tilde{Z}$ be a vertex set maximizing $|\tilde{Z}_{\mathrm{out}}|$ subject to $2|\tilde{Z}_{\mathrm{out}}|>|\Gamma_{\tilde{H}}(\tilde{Z}_{\mathrm{out}})|$, and let $\tilde{Z}_{\mathrm{in}}\coloneqq\tilde{Z}\setminus\tilde{Z}_{\mathrm{out}}$. Then, by the maximality of $\tilde{Z}_{\mathrm{out}}$, we have $2|\tilde{Z}^{\prime}|\leq|\Gamma_{\tilde{H}}(\tilde{Z}^{\prime})|$ for any $\tilde{Z}^{\prime}\subseteq\tilde{Z}_{\mathrm{in}}$. Let $\tilde{Y}_{\mathrm{out}}\coloneqq\Gamma_{\tilde{H}}(\tilde{Z}_{\mathrm{out}})\setminus\Gamma_{\tilde{H}}(\tilde{Z}_{\mathrm{in}})$, $\tilde{Y}_{\mathrm{in}}\coloneqq\tilde{Y}\setminus\tilde{Y}_{\mathrm{out}}$, and $\tilde{H}_{\mathrm{in}}\coloneqq\tilde{H}[\tilde{Z}_{\mathrm{in}}\cup\tilde{Y}_{\mathrm{in}}]$. Then, as $\Gamma_{\tilde{H}}(\tilde{Z}_{\mathrm{in}})\subseteq\tilde{Y}_{\mathrm{in}}$, we have $2|\tilde{Z}^{\prime}|\leq|\Gamma_{\tilde{H}}(\tilde{Z}^{\prime})|=|\Gamma_{\tilde{H}_{\mathrm{in}}}(\tilde{Z}^{\prime})|$ for any $\tilde{Z}^{\prime}\subseteq\tilde{Z}_{\mathrm{in}}$. By Hall’s theorem [16], $\tilde{H}_{\mathrm{in}}$ has two disjoint matchings such that each of them covers $\tilde{Z}_{\mathrm{in}}$ and no vertex in $\tilde{Y}_{\mathrm{in}}$ is covered by both of them. In other words, $\tilde{H}_{\mathrm{in}}$ contains a subgraph $\tilde{H}^{\prime}$ such that $\deg_{\tilde{H}^{\prime}}(\tilde{z})=2$ for every $\tilde{z}\in\tilde{Z}_{\mathrm{in}}$ and $\deg_{\tilde{H}^{\prime}}(\tilde{y})\leq 1$ for every $\tilde{y}\in\tilde{Y}_{\mathrm{in}}$. Let $\tilde{M}\coloneqq E(\tilde{H}^{\prime})$, and $\tilde{Z}(\tilde{M})$ and $\tilde{Y}(\tilde{M})$ be the sets of end vertices of edges in $\tilde{M}$ that are in $\tilde{Z}$ and in $\tilde{Y}$, respectively.
If $\tilde{Z}_{\mathrm{in}}=\emptyset$, then $\tilde{Y}_{\mathrm{out}}=\tilde{Y}$. Hence, $|\tilde{Y}|=|\tilde{Y}_{\mathrm{out}}|<2|\tilde{Z}_{\mathrm{out}}|\leq 2|\tilde{Z}|\leq 2(2k)^{2}=8k^{2}$ (recall that $|X|\leq 4k$ and each vertex $\tilde{z}\in\tilde{Z}$ is a pair of vertices in $G$ one of which is in $X\cap V^{+}$ and the other is in $X\cap V^{-}$). This implies $|V|=|X|+|Y|=|X|+2|\tilde{Y}|\leq 4k(4k+1)$. In this case, the input size is bounded by a function of the parameter $k$, and hence we can solve the problem by an exhaustive search in FPT time.
Otherwise, i.e., if $\tilde{Z}_{\mathrm{in}}\neq\emptyset$, the following claim holds, which is analogous to [2, Claim in p. 2927]. We define $\tilde{Y}^{\prime}\coloneqq\tilde{Y}_{\mathrm{in}}\setminus\tilde{Y}(\tilde{M})$ and $Y^{\prime}\coloneqq\bigcup_{\tilde{y}\in\tilde{Y}^{\prime}}\tilde{y}\ (\subseteq V)$.
Claim 4.5.
There exists a minimum DM-irreducible spanning subgraph $G^{*}$ of $G$ satisfying the following two conditions.
  • $G^{*}-Y^{\prime}$ is a DM-irreducible spanning subgraph of $G-Y^{\prime}$.
  • For any $\tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y}^{\prime}$, there exists $\tilde{z}=\{z^{+},z^{-}\}\in\tilde{Z}$ with $\{\tilde{z},\tilde{y}\}\in E(\tilde{H})$ such that the edges in $E(G^{*})$ incident to $y^{+}$ or $y^{-}$ are $\{z^{+},y^{-}\}$, $\{y^{+},y^{-}\}$, and $\{y^{+},z^{-}\}$.
Before proving Claim 4.5, we complete our FPT algorithm for UnweightedDMISS. By Claim 4.5, the answer is yes if and only if there exists a DM-irreducible spanning subgraph of $G-Y^{\prime}$ consisting of at most $3n-2-k-3|\tilde{Y}^{\prime}|$ edges. We can check all possible spanning subgraphs of $G-Y^{\prime}$ by an exhaustive search in FPT time, since
$|V(G-Y^{\prime})|=|X|+2|\tilde{Y}(\tilde{M})|+2|\tilde{Y}_{\mathrm{out}}|<|X|+4|\tilde{Z}(\tilde{M})|+4|\tilde{Z}_{\mathrm{out}}|=|X|+4|\tilde{Z}|\leq 4k(4k+1).$
Proof of Claim 4.5.
Let $G^{\prime}$ be a minimum DM-irreducible spanning subgraph of $G$. If $G^{\prime}-Y^{\prime}$ is DM-irreducible, then we are done as follows. By Claim 4.4, any DM-irreducible spanning subgraph of $G$ (including $G^{\prime}$) contains, for each pair $\tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y}$, either
  • a path of length $3$ between two vertices in $V(G^{\prime}-Y)$ through the edge $\{y^{+},y^{-}\}$, or
  • at least two edges incident to $y^{+}$ and to $y^{-}$ other than $\{y^{+},y^{-}\}$ (at least four in total).
Thus, a spanning subgraph of $G$ obtained from $G^{\prime}-Y^{\prime}$ as follows is DM-irreducible and has as few edges as $G^{\prime}$ itself, and is a desired graph: for each pair $\tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y}^{\prime}$, choose an arbitrary pair $\tilde{z}=\{z^{+},z^{-}\}\in\tilde{Z}$ with $\{\tilde{z},\tilde{y}\}\in E(\tilde{H})$ and add three edges $\{z^{+},y^{-}\}$, $\{y^{+},y^{-}\}$, and $\{y^{+},z^{-}\}$.
Otherwise, i.e., if $G^{\prime}-Y^{\prime}$ is not DM-irreducible, we transform $G^{\prime}$ as follows: remove all the edges incident to $Y(\tilde{M})\coloneqq\bigcup_{\tilde{y}\in\tilde{Y}(\tilde{M})}\tilde{y}\ (\subseteq V)$, and for each $\{\tilde{z},\tilde{y}\}\in\tilde{M}$, add three edges $\{z^{+},y^{-}\}$, $\{y^{+},y^{-}\}$, and $\{y^{+},z^{-}\}$. Let $G^{\sharp}$ be the resulting graph. We then have $|E(G^{\sharp})|\leq|E(G^{\prime})|$, and it suffices to show that $G^{\sharp}-Y^{\prime}$ is DM-irreducible.
Let $M^{\prime}$ be a perfect matching in $G^{\prime}$. As $G^{\prime}$ is DM-irreducible, $G^{\prime}_{M^{\prime}}$ is strongly connected. We first show that $G^{\sharp}-Y^{\prime}$ has a perfect matching, which we construct from $M^{\prime}$ minus the edges incident to $Y_{\mathrm{in}}\coloneqq\bigcup_{\tilde{y}\in\tilde{Y}_{\mathrm{in}}}\tilde{y}\ (\subseteq V)$ as follows. For each removed edge $\{y^{+},y^{-}\}\in E(G[Y(\tilde{M})])\cap M^{\prime}$, we just add it again. The other removed edges in $M^{\prime}$ are paired as $\{y^{+},z^{-}\}$ and $\{z^{+},y^{-}\}$ such that $\tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y}_{\mathrm{in}}$, $\tilde{z}=\{z^{+},z^{-}\}\in\tilde{Z}_{\mathrm{in}}$, and $\{\tilde{z},\tilde{y}\}\in E(\tilde{H}_{\mathrm{in}})$. By definition of $\tilde{M}$, there exists a pair $\tilde{y}^{\prime}=\{y^{\prime+},y^{\prime-}\}\in\tilde{Y}(\tilde{M})$ with $\{\tilde{z},\tilde{y}^{\prime}\}\in\tilde{M}\subseteq E(\tilde{H}_{\mathrm{in}})$, which means that $\{z^{+},y^{\prime-}\},\{y^{\prime+},y^{\prime-}\},\{y^{\prime+},z^{-}\}\in E(G^{\sharp})$. By definition of $\tilde{M}$ again, such $\tilde{y}^{\prime}$ is distinct for each $\tilde{z}\in\tilde{Z}_{\mathrm{in}}=\tilde{Z}(\tilde{M})$, and hence we can add $\{z^{+},y^{\prime-}\}$ and $\{y^{\prime+},z^{-}\}$ instead of each pair of $\{y^{+},z^{-}\}$ and $\{z^{+},y^{-}\}$. Finally, for each pair $\{y^{+},y^{-}\}\in\tilde{Y}(\tilde{M})$, if at least one of $y^{+}$ and $y^{-}$ is not matched, then both $y^{+}$ and $y^{-}$ are not matched due to the above procedure, and add an edge $\{y^{+},y^{-}\}$. Let $M^{\sharp}$ be the obtained perfect matching in $G^{\sharp}$.
The remaining task is to show that $G^{\sharp}_{M^{\sharp}}-Y^{\prime}$ is strongly connected. By the above construction, it suffices to show that for each pair of two vertices $s,t\in V\setminus Y_{\mathrm{in}}$, $t$ is reachable from $s$ in $G^{\sharp}_{M^{\sharp}}-Y^{\prime}$. Suppose to the contrary that there exists a pair $(s,t)$ such that $t$ is not reachable from $s$ in $G^{\sharp}_{M^{\sharp}}-Y^{\prime}$. As $G^{\prime}_{M^{\prime}}$ is strongly connected, there exists a path $P$ from $s$ to $t$ in $G^{\prime}_{M^{\prime}}$. Take such a pair of $(s,t)$ and $P$ so that the length of $P$ is minimized. Then, all the inner vertices of $P$ are in $Y_{\mathrm{in}}$ (otherwise, we can take an inner vertex $v\not\in Y_{\mathrm{in}}$ such that a pair of $(s,v)$ and the prefix of $P$ or a pair of $(v,t)$ and the suffix of $P$ is eligible, which contradicts that the length of $P$ is minimized), and hence $s,t\in Z_{\mathrm{in}}$.
If $\{s,t\}\in\tilde{Z}_{\mathrm{in}}=\tilde{Z}(\tilde{M})$, then by the above construction, there exists a path from $s$ to $t$ through a pair $\tilde{y}\in\tilde{Y}(\tilde{M})$ in $G^{\sharp}_{M^{\sharp}}-Y^{\prime}$, a contradiction. Otherwise, since $E(G[Y])$ is a matching (Claim 4.4), $P$ consists of two edges $(s,v)$ and $(v,t)$ such that $v\in Y_{\mathrm{in}}$ and exactly one of $\{s,v\}$ and $\{t,v\}$ is in $M^{\prime}$. By symmetry, assume that $s,t\in V^{+}$, $v\in V^{-}$, and $\{t,v\}\in M^{\prime}$. Let $\tilde{y}=\{y^{+},y^{-}\}\in\tilde{Y}_{\mathrm{in}}$ be the pair with $y^{-}=v$. Then, there exist two pairs $\tilde{z}_{1}=\{z_{1}^{+},z_{1}^{-}\}$ and $\tilde{z}_{2}=\{z_{2}^{+},z_{2}^{-}\}$ in $\tilde{Z}_{\mathrm{in}}=\tilde{Z}(\tilde{M})$ such that $z_{1}^{+}=s$, $z_{2}^{+}=t$, $z_{1}^{-}=z_{2}^{-}$ and $\{y^{+},z_{1}^{-}\}\in M^{\prime}$. By the above construction, there exist two pairs $\tilde{y}_{1}=\{y_{1}^{+},y_{1}^{-}\}$ and $\tilde{y}_{2}=\{y_{2}^{+},y_{2}^{-}\}$ in $\tilde{Y}(\tilde{M})$ such that $\{\tilde{z}_{1},\tilde{y}_{1}\},\{\tilde{z}_{2},\tilde{y}_{2}\}\in\tilde{M}$ and $\{y_{1}^{+},y_{1}^{-}\},\{z_{2}^{+},y_{2}^{-}\},\{y_{2}^{+},z_{2}^{-}\}\in M^{\sharp}$. This concludes that $G^{\sharp}_{M^{\sharp}}-Y^{\prime}$ has a path from $s=z_{1}^{+}$ to $t=z_{2}^{+}$ that consists of six edges $(s,y_{1}^{-})$, $(y_{1}^{-},y_{1}^{+})$, $(y_{1}^{+},z_{1}^{-})$, $(z_{2}^{-},y_{2}^{+})$, $(y_{2}^{+},y_{2}^{-})$, and $(y_{2}^{-},t)$, a contradiction. Thus, we are done. ∎
Remark 4.6.
Another possible parameterized approach is, as a connectivity augmentation problem (like [3]), taking the parameter $k$ as the number of additional edges to achieve DM-irreducibility. For strong connectivity augmentation, Klinkby, Misra, and Saurabh [20] recently showed the first FPT algorithm in this direction. How to extend their result to the DM-irreducibility problem is highly nontrivial because there are too many possible choices of perfect matchings in the resulting augmented graphs. This may be possible future work.

### v3

@@
 Abstract
-Finding a minimum-weight strongly connected spanning subgraph of an edge-weighted directed graph is equivalent to the weighted version of the well-known strong connectivity augmentation problem, and it is NP-hard.
-For this problem, a simple $2$-approximation algorithm was proposed by Frederickson and Jájá (1981); surprisingly, it still achieves the best-known approximation ratio in general.
-Also, the unweighted problem was shown to be FPT (fixed-parameter tractable) by Bang-Jensen and Yeo (2008), if it is parameterized by the difference from the trivial upper bound of the optimal value.
-In this paper, we consider a generalized problem related to the Dulmage–Mendelsohn decompositions of bipartite graphs instead of the strong connectivity of directed graphs, and extend these approximation and FPT results to the generalized setting.
+Finding a minimum-weight strongly connected spanning subgraph of an edge-weighted directed graph is equivalent to the weighted version of the well-known strong connectivity augmentation problem.
+This problem is NP-hard, and a simple $2$-approximation algorithm was proposed by Frederickson and Jájá (1981); surprisingly, it still achieves the best known approximation ratio in general.
+Also, Bang-Jensen and Yeo (2008) showed that the unweighted problem is FPT (fixed-parameter tractable) parameterized by the difference from a trivial upper bound of the optimal value.
+In this paper, we consider a generalization related to the Dulmage–Mendelsohn decompositions of bipartite graphs instead of the strong connectivity of directed graphs, and extend these approximation and FPT results to the generalized setting.
 Keywords
@@
 The strong connectivity augmentation problem (SCA) is, given a directed graph and the weights of additional edges, to find a minimum-weight subset of additional edges making the graph strongly connected.
-While SCA is NP-hard in general, it can be solved in linear time in the unweighted case (when any edge can be added and has the same weight) [9].
+While SCA is NP-hard in general, it can be solved in linear time in the unweighted case (when an edge between any pair of veritces can be added with the same weight) [10].
 The following problem, called the strongly connected spanning subgraph problem (SCSS), is polynomial-time equivalent to SCA in general: given a strongly connected directed graph with edge weight, find a minimum-weight spanning subgraph that is strongly connected.
+This problem is one of the most fundamental special cases of the survivable network design problem [27, 6].
 Frederickson and Jájá [15] gave a $2$-approximation algorithm for SCSS; this is based on a very simple observation (see Section 3.1) and, surprisingly, it still achieves the best known approximation ratio in general.
 For the unweighted case (when all the edges have the same weight, which is already NP-hard in contrast to SCA), several algorithms with better approximation ratios [19, 28, 29, 18] have been proposed.
-Also, Bang-Jensen and Yeo [2] showed that the unweighted case is FPT if it is parameterized by the difference from the trivial upper bound of the optimal value (see Sections 2.2 and 4.1); here, the problem is called FPT (fixed-parameter tractable) if it can be solved in $f(k)\cdot n^{O(1)}$ time, where $k$ and $f$ are the prescribed parameter and computable function, respectively, and $n$ is the input size.
+Also, Bang-Jensen and Yeo [2] showed that the unweighted case is FPT parameterized by the difference from a trivial upper bound of the optimal value (see Sections 2.2 and 4.1); here, the problem is called FPT (fixed-parameter tractable) if it can be solved in $f(k)\cdot n^{O(1)}$ time, where $k$ and $f$ are the prescribed parameter and computable function, respectively, and $n$ is the input size.
 SCSS (or SCA) is generalized to several directions.
-One natural generalization is the $k$(-edge)-connected spanning subgraph (or $k$(-edge)-connectivity augmentation) problem, where a directed graph is called $k$-connected (or $k$-edge-connected) if there exists $k$ openly disjoint (or edge-disjoint) paths between any pair of vertices.
+One natural generalization is the $k$(-edge)-connected spanning subgraph (or $k$(-edge)-connectivity augmentation) problem, where a directed graph is called $k$-connected (or $k$-edge-connected) if there exists $k$ openly vertex-disjoint (or edge-disjoint) paths between any pair of vertices.
 The case when $k=1$ corresponds to SCSS (or SCA), and the general case is also well-studied [11, 13, 20, 25, 1].
-Bérczi, Iwata, Kato, and Yamaguchi [3] recently investigated another generalization of SCA, which is related to the Dulmage–Mendelsohn decomposition (the DM-decomposition) [7, 6] of a bipartite graph.
+Bérczi, Iwata, Kato, and Yamaguchi [3] recently introduced another generalization of SCA, which is related to the Dulmage–Mendelsohn decomposition (the DM-decomposition) [8, 7] of a bipartite graph.
 The DM-decomposition gives a partition of the vertex set that reflects the structure of all maximum matchings therein, and a bipartite graph is DM-irreducible if its DM-decomposition consists of a single component.
 This property is closed under adding edges, and we analogously formulate the DM-irreducible spanning subgraph problem (DMISS): given a DM-irreducible bipartite graph with edge weight, find a minimum-weight spanning subgraph that is DM-irreducible.
-Then, DMISS indeed generalizes SCSS (cf.
-[3, Section 2.3]), and in terms of DMISS, Bérczi, et al.
-[3] proposed polynomial-time algorithms for the case when the input bipartite graph is unbalanced with the aid of weighted matroid intersection, and for the case when the input graph is a balanced complete bipartite graph and the weight of each edge is 00 or $1$, where a bipartite graph is balanced if the two color classes have the same number of vertices.
-In this paper, by extending the approaches to SCSS by Frederickson and Jájá [14] and by Bang-Jensen and Yeo [2], we propose a $2$-approximation algorithm for DMISS in general and an FPT algorithm for the unweighted case of DMISS.
+In terms of DMISS, Bérczi et al.
+[3] proposed polynomial-time algorithms for the case when the input bipartite graph is unbalanced (which reduces to the weighted matroid intersection problem), and for the case when the input graph is a balanced complete bipartite graph and the weight of each edge is 00 or $1$; here a bipartite graph is balanced if the two color classes have the same number of vertices.
+DMISS for balanced bipartite graphs generalizes SCSS (cf.
+[3, Section 2.3]), and the latter result extends the polynomial-time solvability of the unweighted case of SCA [10].
+In this paper, we investigate DMISS for balanced bipartite graphs, whose algorithmic difficulty has remained widely open (other than the fact that it is NP-hard even in the unweighted case as so is SCSS).
+As natural but nontrivial results, by extending the approaches to SCSS by Frederickson and Jájá [15] and by Bang-Jensen and Yeo [2], we propose a $2$-approximation algorithm in general and an FPT algorithm for the unweighted case.
+On the approximation result, we first show that a naive extension of [15] gives a $3$-approximation algorithm, and then improve it to a $2$-approximation algorithm with the aid of the matroid intersection framework.
+On the FPT result, we demonstrate that the idea of [2] can be extended by carefully adjusting several key concepts and discussions.
 The rest of this paper is organized as follows.
@@
 In Section 3, we give a $2$-approximation algorithm for DMISS.
-In Section 4, we prove the unweighted case of DMISS is FPT parameterized by the difference from the trivial upper bound of the optimal value.
+In Section 4, we prove the unweighted case of DMISS is FPT parameterized by the difference from a trivial upper bound of the optimal value.
 2 Preliminaries
@@
 A subgraph $G^{\prime}$ of $G$ is induced by a vertex set $X\subseteq V(G)$, denoted by $G[X]$, if $V(G^{\prime})=X$ and $E(G^{\prime})$ is the set of edges in $E(G)$ whose end vertices are both in $X$.
-We also define $G-X\coloneqq G[V\setminus X]$ for $X\subseteq V$.
+We also define $G-X\coloneqq G[V(G)\setminus X]$ for $X\subseteq V(G)$.
 A subgraph $G^{\prime}$ of $G$ is spanning if $V(G^{\prime})=V(G)$.
 We often do not distinguish between a spanning subgraph $G^{\prime}$ and its edge set $E(G^{\prime})$.
-A stable set in an undirected graph $G$ is a vertex set $X\subseteq V(G)$ such that no edge in $E(G)$ has its both end vertices in $X$.
-An undirected graph $G=(V,E)$ is bipartite if there exists a bipartition $(V^{+},V^{-})$ of $V$ such that $V^{+}\cap V^{-}=\emptyset$, $V^{+}\cup V^{-}=V$, and both $V^{+}$ and $V^{-}$ are stable sets in $G$ (possibly $V^{+}$ or $V^{-}$ is empty).
+Also, we often specify a graph $G$ as the pair $(V,E)$ of its vertex set $V=V(G)$ and edge set $E=E(G)$.
+Let $G=(V,E)$ be an undirected graph.
+A stable set in $G$ is a vertex set $X\subseteq V$ such that no edge in $E$ has its both end vertices in $X$.
+We say that $G$ is bipartite if there exists a bipartition $(V^{+},V^{-})$ of $V$ such that $V^{+}\cap V^{-}=\emptyset$, $V^{+}\cup V^{-}=V$, and both $V^{+}$ and $V^{-}$ are stable sets in $G$ (possibly $V^{+}$ or $V^{-}$ is empty).
 To specify the bipartition, we denote by $G=(V^{+},V^{-};E)$.
@@
 The DM-decomposition [8, 7] of a bipartite graph is formally defined by the distributive lattice formed by the minimizers of a submodular function, and a bipartite graph is called DM-irreducible if its DM-decomposition consists of a single component.
-We here employ its characterization as an alternative “definition” (cf.
+We here skip the formal definition and instead employ its characterization as an alternative “definition” (cf.
 [3, Section 2.2]).
 [^1] The equivalent condition is also called elementary, e.g., in [24, 12].
-Lemma 2.1 (cf. [11, Theorem 2.4.18]).
+Lemma 2.1 (cf. [12, Theorem 2.4.18]).
 For a balanced bipartite graph $G$, the following statements are equivalent.
@@
   3. 3.
-    $G$ has a perfect matching, and for any perfect matching $M$ in $G$, $G_{M}$ is strongly connected.
-Suppose that a graph $G$ is associated with edge weight $w\colon E\to\mathbb{R}$. We define the weight of each edge set $F\subseteq E$ as $w(F)\coloneq\sum_{e\in F}w(e)$.
+    $G$ has a perfect matching and $G_{M}$ is strongly connected for any perfect matching $M$ in $G$.
+Suppose that a graph $G$ is associated with edge weight $w\colon E(G)\to\mathbb{R}$. We define the weight of each edge set $F\subseteq E(G)$ as $w(F)\coloneq\sum_{e\in F}w(e)$.
 2.2 Problems and Our Results
-We formally state the problems dealt with in this paper. Recall that DMISS for the unbalanced bipartite graphs is tractable [3], and it is not considered in this paper.
+We formally state the problems dealt with in this paper. Recall that DMISS for unbalanced bipartite graphs is polynomial-time solvable [3], which is not considered here.
 Problem (Strongly Connected Spanning Subgraph (SCSS)).
 Input:
-A strongly connected digraph $G=(V,E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
+A strongly connected digraph $G$ with $w\colon E(G)\to\mathbb{R}_{\geq 0}$.
 Goal:
@@
 Input:
-A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
+A DM-irreducible balanced bipartite graph $G$ with $w\colon E(G)\to\mathbb{R}_{\geq 0}$.
 Goal:
@@
 It is easy to observe that any inclusion-wise minimal strongly connected spanning subgraph of a digraph on $n$ vertices consists of at most $22n-$ 2 edges (by Lemma 3.1).
-Similarly, it is easy to observe that any inclusion-wise minimal DM-irreducible spanning subgraph of a balanced bipartite graph on $2n$ vertices ($n$ vertices on each side) consists of at most $23n-$ 2 edges (by Lemma 3.2).
-The following problems are the unweighted version of the above two problems parameterized by the difference from these trivial upper bounds (note that it is obvious from the definition that $|V|$ edges are necessary for either problem).
+Similarly, any inclusion-wise minimal DM-irreducible spanning subgraph of a balanced bipartite graph on $2n$ vertices ($n$ vertices on each side) consists of at most $23n-$ 2 edges (by Lemma 3.2).
+The following problems are the unweighted version of the above two problems parameterized by the difference from these trivial upper bounds; note that it is obvious from the definition that $|V(G)|$ edges are necessary in either problem.
 Problem (UnweightedSCSS).
 Input:
-A strongly connected digraph $G=(V,E)$.
+A strongly connected digraph $G$ on $n$ vertices.
 Parameter:
-An integer $k\leq|V|-2$.
+An integer $k\leq n-2$.
 Goal:
-Test whether $G$ has a strongly connected spanning subgraph with at most $2|V|-2-k$ edges or not.
+Test whether $G$ has a strongly connected spanning subgraph with at most $2n-2-k$ edges or not.
 Problem (UnweightedDMISS).
 Input:
-A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$.
+A DM-irreducible balanced bipartite graph $G$ on $2n$ vertices.
 Parameter:
-An integer $k\leq|V^{+}|-2$.
+An integer $k\leq n-2$.
 Goal:
-Test whether $G$ has a DM-irreducible spanning subgraph with at most $3|V^{+}|-2-k$ edges or not.
-The results shown in this paper are summarized as follows, where $n=|V^{+}|=|V^{-}|$.
+Test whether $G$ has a DM-irreducible spanning subgraph with at most $3n-2-k$ edges or not.
+The results of this paper are summarized as follows, where $n=\frac{1}{2}|V(G)|$.
 Theorem 2.2.
@@
 In this section, we prove Theorem 2.2 by presenting a $2$-approximation algorithm for DMISS that runs in $O(n^{3})$ time.
-In Section 3.1, as a base of our result, we describe a $2$-approximation algorithm for SCSS by Frederickson and Jájá [14].
+In Section 3.1, as a base of our result, we review the $2$-approximation algorithm for SCSS proposed by Frederickson and Jájá [15].
 In Section 3.2, we show a $3$-approximation algorithm for DMISS obtained by a naive extension of the result of Frederickson and Jájá.
@@
 3.1 2-approximation Algorithm for SCSS
-The algorithm is based on the following characterization of strongly connected digraphs.
+We start with an elementary characterization of strongly connected digraphs.
 Lemma 3.1.
@@
     For any vertex $r$, $G$ has a spanning in-arborescence and out-arborescence rooted at $r$.
-For the input $(G=(V,E),w)$ of SCSS, the algorithm arbitrarily chooses a vertex $r\in V$ and finds a minimum-weight spanning in-arborescence $A_{\mathrm{in}}$ and out-arborescence $A_{\mathrm{out}}$ rooted at $r$; the existence of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ follows from Lemma 3.1 $(1\Rightarrow 3)$.
-This can be done in $O(m+n\log n)$ time [15], where $n=|V|$ and $m=|E|$.
+For the input $(G,w)$ of SCSS, the algorithm arbitrarily chooses a vertex $r$ and finds a minimum-weight spanning in-arborescence $A_{\mathrm{in}}$ and out-arborescence $A_{\mathrm{out}}$ rooted at $r$; the existence of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ follows from Lemma 3.1 $(1\Rightarrow 3)$.
+This can be done in $O(m+n\log n)$ time [16], where $n=|V(G)|$ and $m=|E(G)|$.
 By Lemma 3.1 $(2\Rightarrow 1)$, the union of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$ results in a strongly connected spanning subgraph of $G$, and the algorithm outputs this.
-Since any optimal solution has a spanning in-arborescence and out-arborescence rooted at $r$ by Lemma 3.1 $(1\Rightarrow 3)$ and their weights are at least those of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$, respectively, the weight of the output solution is at most twice the optimal value.
+Since an optimal solution has a spanning in-arborescence and out-arborescence rooted at $r$ by Lemma 3.1 $(1\Rightarrow 3)$ and their weights are at least those of $A_{\mathrm{in}}$ and $A_{\mathrm{out}}$, respectively, the weight of the output solution is at most twice the optimal value.
 3.2 3-approximation Algorithm for DMISS
@@
 We then output the union of the solutions (for the second solution, we take the set of corresponding edges in $G$).
-This intuitively leads to a $3$-approximation algorithm for DMISS (as $1+2=3$), which can be formally shown as follows.
+This intuitively leads to a $3$-approximation algorithm for DMISS, which can be formally shown as follows.
 Lemma 3.3.
@@
 We improve the above $3$-approximation algorithm with the aid of matroid intersection.
-We first give an easy but important observation. In short, for any $r\in V^{+}$, all the matching edges are included in every spanning $r$-in-arborescence in the auxiliary graph.
+We first give an easy but important observation.
 Lemma 3.4.
@@
   1. 1.
-    $T$ has a perfect matching $M$ in $G$ and $\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is a spanning in-arborescence of $G_{M}$.
+    $T$ has a perfect matching $M$ in $G$ and $\overrightarrow{T\setminus M}\cup\overleftarrow{M}$ is a spanning in-arborescence of $G_{M}$ (rooted at some vertex in $V^{+}$).
   2. 2.
@@
 Let us call a spanning tree satisfying Statement 2 in Lemma 3.5 strongly balanced, and say that the vertex $r$ is the root and the tree is rooted at $r$.
-By Lemma 3.2, every DM-irreducible graph has a strongly balanced spanning tree, and then we formulate the first task as follows.
+By Lemma 3.2, every DM-irreducible balanced bipartite graph has a strongly balanced spanning tree, and then we formulate the first task as follows.
 Problem (Strongly Balanced Spanning Tree (SBST)).
 Input:
-A DM-irreducible bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$.
+A DM-irreducible balanced bipartite graph $G$ with $w\colon E(G)\to\mathbb{R}_{\geq 0}$.
 Goal:
 Find a minimum-weight strongly balanced spanning tree of $G$.
-For the sake of simplicity of the following discussion, we consider the following problem, which is polynomially equivalent to the above problem and actually enough for our purpose.
-Problem (Strongly Balanced Spanning Tree with Root (SBSTR)).
+For the sake of simplicity of the following discussion, we consider the following problem, which is polynomial-time equivalent to the above problem and actually enough for our purpose.
+Problem (SBST with Root (SBSTR)).
 Input:
-A DM-irreducible bipartite graph $G=(V^{+},V^{-};E)$ with edge weight $w\colon E\to\mathbb{R}_{\geq 0}$, and $r\in V^{+}$.
+A DM-irreducible balanced bipartite graph $G=(V^{+},V^{-};E)$ with $w\colon E\to\mathbb{R}_{\geq 0}$, and $r\in V^{+}$.
 Goal:
@@
 Computational Time
-As we already mentioned, the second task (finding a minimum-weight spanning arborescence) can be done in $O(m+n\log n)$ time [15].
+Let $n=|V^{+}|=|V^{-}|$ and $m=|E|$. As we already mentioned, the second task (finding a minimum-weight spanning arborescence) can be done in $O(m+n\log n)$ time [16].
 The following observation (which was also observed in [14]) leads to an $O(n^{3})$-time algorithm for the first task (solving SBSTR), which completes the analysis.
 We omit the basics on matroids, for which we refer the readers to [26].
-Lemma 3.6 (cf. [13, Lemma 5]).
+Lemma 3.6 (cf. [14, Lemma 5]).
 SBST and SBSTR are formulated as the weighted matroid intersection problem, i.e., the set of feasible solutions can be represented as the set of common bases of two matroids.
@@
 with the aid of matroid intersection algorithms.
-In particular, Brezovec, Cornuéjols, and Glover [5] proposed an $O(\tilde{n}\tilde{m}+\tilde{k}\tilde{n}^{2})$-time algorithm for the following special case: one matroid is a cycle matroid of a graph and the other is a partition matroid whose partition classes are induced by a stable set of the graph (i.e., each partition class is the set of edges incident to a vertex in the stable set), where $\tilde{n}$ and $\tilde{m}$ are the numbers of vertices and of edges in the graph, respectively, and $k$ is the size of the stable set inducing the partition classes.
-By the proof of Lemma 3.6, SBSTR completely fits this situation, and hence we can exactly solve it in $O(n^{3})$ time, where we have $\tilde{k}=|V^{+}|=n=\frac{\tilde{n}}{2}$ and $\tilde{m}\leq n^{2}$.
+In particular, Brezovec, Cornuéjols, and Glover [5] proposed an $O(\tilde{n}\tilde{m}+\tilde{k}\tilde{n}^{2})$-time algorithm for the following special case: one matroid is the cycle matroid of a graph and the other is a partition matroid whose partition classes are induced by a stable set of the graph (i.e., each partition class is the set of edges incident to a vertex in the stable set), where $\tilde{n}$ and $\tilde{m}$ are the numbers of vertices and of edges in the graph, respectively, and $\tilde{k}$ is the size of the stable set inducing the partition classes.
+By the proof of Lemma 3.6, SBSTR completely fits this situation, and hence we can exactly solve it in $O(n^{3})$ time, where we have $\tilde{k}=n=\frac{\tilde{n}}{2}$ and $\tilde{m}\leq n^{2}$.
 Remark 3.7.
@@
 Let $S_{\mathrm{opt}}$ be an optimal solution.
-We show $w(T\cup\overline{A_{\mathrm{out}}})=w(T)+w_{M}(A_{\mathrm{out}})\leq 2w(S_{\mathrm{opt}})$ (recall that $w_{M}\colon E(G_{M})\to\mathbb{R}_{\geq 0}$ is defined by (1)).
+We show $w(T\cup\overline{A_{\mathrm{out}}})=w(T)+w_{M}(A_{\mathrm{out}})\leq 2w(S_{\mathrm{opt}})$ (recall that $w_{M}\colon E(G_{M})\to\mathbb{R}_{\geq 0}$ is defined as (1)).
 The following two claims imply this.
@@
 4.1 FPT Algorithm for Unweighted SCSS
-The algorithm utilizes another well-known characterization of strongly connected digraphs.
+The FPT algorithm utilizes another well-known characterization of strongly connected digraphs.
 For a digraph $G=(V,E)$, an ear decomposition of $G$ is a sequence $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of edge-disjoint paths and cycles in $G$ with $E=\bigcup_{i=0}^{f}E(P_{i})$ such that
   • $P_{0}$ is a cycle, and
-  • $P_{i}$ is either a path between two different vertices in $\bigcup_{j=0}^{i-1}V(P_{j})$ whose inner vertices are in $V\setminus\bigcup_{j=0}^{i-1}V(P_{j})$, or a cycle intersecting exactly one vertex in $\bigcup_{j=0}^{i-1}V(P_{j})$.
-Each path or cycle $P_{i}$ is called an ear. An ear is called long if it is of length at least $3$.
-Lemma 4.1 (cf. [11, Theorem 2.2.1]).
+  • $P_{i}$ $(i=1,2,\dots,f)$ is either a path between two different vertices in $\bigcup_{j=0}^{i-1}V(P_{j})$ whose inner vertices are in $V\setminus\bigcup_{j=0}^{i-1}V(P_{j})$, or a cycle intersecting exactly one vertex in $\bigcup_{j=0}^{i-1}V(P_{j})$.
+Each path or cycle $P_{i}$ is called an ear. An ear is called long if it is of length at least $3$, where the length is evaluated by the number of edges.
+Lemma 4.1 (cf. [12, Theorem 2.2.1]).
 A digraph is strongly connected if and only if it has an ear decomposition.
-The idea of the FPT algorithm for UnweightedSCSS in [2] is sketched as follows.
-It first computes an ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of the input graph $G$ so that the set of long ears is inclusion-wise maximal (i.e., so that no more long ear can be taken in addition to them, but the number of long ears in $\mathcal{E}$ is not necessarily maximized).
+The idea of the FPT algorithm for UnweightedSCSS is sketched as follows.
+It first computes an ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ of the input graph $G$ so that the set of long ears is inclusion-wise maximal (i.e., the set of long ears in $\mathcal{E}$ is not a proper subset of that in any ear decomposition $\mathcal{E}^{\prime}$ of $G$, but the number of long ears in $\mathcal{E}$ is not necessarily maximized).
 If the number of long ears is large enough (depending on the parameter $k$), then the ear decomposition $\mathcal{E}$ gives a strongly connected spanning subgraph that is small enough by removing all the ears of length $1$, and the answer is yes.
@@
 Under the assumption that $\mathcal{E}$ does not immediately give a solution, we can show that $|X|=O(k)$.
-Also, by the maximality of long ears in $\mathcal{E}$, the vertices in $V\setminus X$ are added by ears of length $2$, and this is basically true for some minimum strongly connected spanning subgraph.
+Also, by the inclusion-wise maximality of long ears in $\mathcal{E}$, the vertices in $V\setminus X$ are added by ears of length $2$, and this is basically true for some minimum strongly connected spanning subgraph.
 Thus, we can test the existence of a solution by an exhaustive search on the subgraphs of $G$ on the vertex set $X$ (plus a small subset), where we have only $2^{O(k^{2})}$ candidates.
@@
 First, we give an alternative characterization of DM-irreducible balanced bipartite graphs by ear decompositions.
-An ear decomposition of a bipartite graph $G=(V^{+},V^{-};E)$ is defined in the same way as above, and we call it odd proper if
+An ear decomposition of a bipartite graph $G=(V^{+},V^{-};E)$ is defined in the same way, and we call it odd proper if
   • $P_{0}$ is a cycle of length at least $4$, and
   • each $P_{i}$ $(i=1,2,\dots,f)$ is a path of odd length, which implies that one end vertex of $P_{i}$ is in $V^{+}$ and the other is in $V^{-}$.
-Lemma 4.2 (cf. [11, Theorem 2.4.17] or [23, Theorem 4.1.6]).
+Lemma 4.2 (cf. [12, Theorem 2.4.17] or [24, Theorem 4.1.6]).
 A bipartite graph on at least four vertices is balanced and DM-irreducible if and only if it has an odd proper ear decomposition.
@@
 [^3] This redefinition is reasonable by considering the relation between the strong connectivity of a digraph and the DM-irreducibility of a balanced bipartite graph (cf.
-[3, Section 2.3])
+[3, Section 2.3]); a path of length $\ell$ in the former corresponds to a path of length $12\ell-$ 1 in the latter (as a vertex in the former is split into two adjacent vertices in the latter).
 We start to describe our FPT algorithm for UnweightedDMISS.
@@
 We first find an odd proper ear decomposition $\mathcal{E}=(P_{0},P_{1},\dots,P_{f})$ such that
-  • $\{P_{1},P_{2},\dots,P_{s}\}$ is the set of long ears in $\mathcal{E}$, which is inclusion-wise maximal (i.e., no more long ear can be taken in addition to them, where $s$ is not necessarily maximized), and
+  • $\{P_{1},P_{2},\dots,P_{s}\}$ is the set of long ears in $\mathcal{E}$, which is inclusion-wise maximal (i.e., the set of long ears in $\mathcal{E}$ is not a proper subset of that in any ear decomposition $\mathcal{E}^{\prime}$ of $G$, but $s$ is not necessarily maximized), and
   • $\{P_{1},P_{2},\dots,P_{r}\}$ is the set of nontrivial ears in $\mathcal{E}$ (note that $s\leq r\leq f$).
@@
 Claim 4.4.
-$E(G[Y])$ is a perfect matching.
+$E(G[Y])$ is a perfect matching in $G[Y]$.
 Proof.